# Learning Regularization for Graph Inverse Problems

### v1

@@ Appendix G Complexity and Runtimes
To make a quantitative comparison of the runtimes for the various trainable models, we measured the training and inference times for each. Namely, we measure the runtimes on the CLUSTER (for the graph property completion task) and CPOX (for the inverse source estimation task) for $nb=4$ and $k=4$ respectively. The GPU runtimes were measured in milliseconds per batch for training and inference. The results are provided in Table 14, which includes the number of trainable parameters for each model. While solutions like Var-GNN and ISS-GNN require more time and parameters than Prox-GNN, they typically offer better performance. Therefore, our framework provides several possible architectures that can be chosen depending on the needs, i.e., better downstream performance or lighter computational footprint. prox-GNN does not require many parameters by design, and so in the table we also provide results with a variant of Prox-GNN that we name Prox-GNN (wide) such that it has a similar parameter count to Var-GNN and ISS-GNN. This was achieved by increasing the number of channels from the CPOX $k=4$ hyper-parameters for Prox-GNN in Table 12 from $64$ to $255$, and by increasing the number of channels from the CLUSTER $nb=4$ hyper-parameters for Prox-GNN (in Table 10) from $128$ to $183$. Note in particular that having a similar number of parameters as the other two models does not lead to Prox-GNN (wide) exhibiting comparable performance to Var-GNN and ISS-GNN.
In terms of complexity, the framework proposed here involves several iterations (denoted by ’solverIter’) of the backbone GNN and data fit steps using conjugate gradient iterations which are of linear complexity (denoted by ’cglsIter’). The GNN backbone used here has a linear complexity with respect to the number of nodes and edges of the graph. Overall, the proposed networks Var-GNN, ISS-GNN, and Prox-GNN maintain the same linear space and time complexity, similar to the backbone GNN.
@@
Table 10: The hyper-parameters for CLUSTER (inv. source estimation) and CLUSTER (prop. completion).
@@
CLUSTER (inv. source estimation) ISS-GNN $k$ = 4 batch_size = 64, lr = 0.00583, wd = 3.43e-05, layers = 8, channels = 32, cglsIter = 32, solveIter = 16
@@
Var-GNN $k$ = 4 batch_size=4, cglsIter=8, channels=64, layers=8, lr=0.0018, solveIter=16, wd=3.81e-06
$k$ = 8 batch_size 8, cglsIter=8, channels=64, layers 8, lr=0.0018, solveIter=16, wd=3.81e-06
$k$ = 16 batch_size 8, cglsIter=8, channels=64, layers 8, lr=0.0018, solveIter=16, wd=3.81e-06
@@
Table 11: The hyper-parameters for ShapeNet (graph property completion).
@@
Table 12: The hyper-parameters for METR-LA (inverse graph transport) and CPOX (inverse source estimation). The number of diffusion steps is denoted by $k$ for CPOX. The path length is denoted by $pl$ for METR-LA
@@
$k$ = 16 batch_size=1, step_size=0.00020, solveIter=3000
@@
Table 16: Inverse Graph Transport normalized-mean-squared-error (MSE) on the METR-LA dataset, showing the nMSE of the recovery loss on ${\bf x}$ and the nMSE of the data fit.
@@
Table 17: Inverse Source Estimation mean-squared error (MSE) of the error and data-fit on the Chickenpox-Hungary dataset.

### v2

@@ Appendix G Complexity and Runtimes
To make a quantitative comparison of the runtimes for the various trainable models, we measured the training and inference times for each. Namely, we measure the runtimes on the CLUSTER (for the graph property completion task) and CPOX (for the inverse source estimation task) for $nb=4$ and $k=4$ respectively. The GPU runtimes were measured in milliseconds per batch for training and inference. The results are provided in Table 14, which includes the number of trainable parameters for each model. While solutions like Var-GNN and ISS-GNN require more time and parameters than Prox-GNN, they typically offer better performance. Therefore, our framework provides several possible architectures that can be chosen depending on the needs, i.e., better downstream performance or lighter computational footprint. prox-GNN does not require many parameters by design, and so in the table, we also provide results with a variant of Prox-GNN that we name Prox-GNN (wide) such that it has a similar parameter count to Var-GNN and ISS-GNN. This was achieved by increasing the number of channels from the CPOX $k=4$ hyper-parameters for Prox-GNN in Table 12 from $64$ to $255$, and by increasing the number of channels from the CLUSTER $nb=4$ hyper-parameters for Prox-GNN (in Table 10) from $128$ to $183$. Note in particular that having a similar number of parameters as the other two models does not lead to Prox-GNN (wide) exhibiting comparable performance to Var-GNN and ISS-GNN.
In terms of complexity, the framework proposed here involves several iterations (denoted by ’solverIter’) of the backbone GNN and data fit steps using conjugate gradient iterations which are of linear complexity (denoted by ’cglsIter’). The GNN backbone used here has a linear complexity with respect to the number of nodes and edges of the graph. Overall, the proposed networks Var-GNN, ISS-GNN, and Prox-GNN maintain the same linear space and time complexity with respect to the number of nodes $n$ and edges $m$ in the graph, similar to the backbone GNNs used in this work, such as GCN (Kipf and Welling 2016).
@@
Table 10: The hyper-parameters for CLUSTER (inverse source estimation) and CLUSTER (property completion).
@@
CLUSTER (inv. source estimation) ISS-GNN $k$ = 4 batch_size = 8, lr = 0.00583, wd = 3.43e-05, layers = 8, channels = 32, cglsIter = 32, solveIter = 16
@@
Var-GNN $k$ = 4 batch_size=8, cglsIter=8, channels=64, layers=8, lr=0.0018, solveIter=16, wd=3.81e-06
$k$ = 8 batch_size=8, cglsIter=8, channels=64, layers 8, lr=0.0018, solveIter=16, wd=3.81e-06
$k$ = 16 batch_size=8, cglsIter=8, channels=64, layers 8, lr=0.0018, solveIter=16, wd=3.81e-06
@@
Table 11: The hyper-parameters for ShapeNet (property completion).
@@
Table 12: The hyper-parameters for METR-LA (inverse graph transport) and CPOX (inverse source estimation). The number of diffusion steps is denoted by $k$ for CPOX. The path length is denoted by $pl$ for METR-LA.
@@
$k$ = 16 batch_size=1, step_size=0.0002, solveIter=3000
@@
Table 16: Inverse Graph Transport normalized-mean-squared-error (nMSE) on the METR-LA dataset, showing the nMSE of the recovery of ${\bf x}$ and the nMSE of the data fit.
@@
Table 17: Inverse Source Estimation normalized-mean-squared error (nMSE) of the recovery of ${\bf x}$ and the data-fit on the Chickenpox-Hungary dataset.
@@
Hyperparameter Changed Accuracy (%) Data Fit (CE)
As in main paper 89.16 0.24
cglsIter=4 87.93 0.51
cglsIter=16 89.14 0.18
channels=128 89.10 0.23
channels=32 88.67 0.26
layers=8 88.54 0.20
layers=4 88.13 0.22
lr=0.001 89.07 0.23
lr=0.01 88.52 0.27
solveIter=16 89.14 0.24
solveIter=4 88.87 0.28
wd=5e-05 89.10 0.25
wd=5e-04 88.94 0.22
Table 19: Comparison of hyperparameter changes with accuracy and data fit (cross-entropy loss).